OpenStudy (anonymous):

Find the solutions to 0 > –3x4 + 5x2 + 1 by graphing f(x) > –3x4 + 5x2 + 1 using technology

OpenStudy (anonymous):

use the zero function on the calculator

OpenStudy (anonymous):

howw please help i only have a scientific calculator i know the solutions are -.3 3.3 3.56 but i dont know how to write it in terms of x

OpenStudy (anonymous):

this technology doesnt exist yet

OpenStudy (anonymous):

yes it does. http://www.wolframalpha.com/input/?i=y+%3E+ –3x%5E4+%2B+5x%5E2+%2B+1 You are looking for where the function is less than 0, so when it is below the y axis. For what values of x does this apply? Well consider the function, where does it cross the axis? http://www.wolframalpha.com/input/?i=0+%3D+ –3x%5E4+%2B+5x%5E2+%2B+1 \[x = \pm \sqrt{\frac{ 5+\sqrt{37} }{ 6 }}\] So the solution is x > √((5+√37)/6) and x < -√((5+√37)/6)